OpenStudy (anonymous):

You and a friend go biking. You bike 12 miles north and 2 miles east. What is the straight-line distance from your starting point? Round answer to the nearest hundredths.

OpenStudy (anonymous):

I would use Pythagoras theorem \[a^2+b^2=c^2\] where a=12 b=2 c=straight-line distance from your starting point So just plug in a b and solve for c

OpenStudy (anonymous):

oh thank you so much

OpenStudy (anonymous):

so it would be 24?

OpenStudy (anonymous):

\[12^2+2^2=c^2\]\[144+4=c^2\]\[148=c^2\] \[c=\sqrt{148}\]

OpenStudy (anonymous):

simplify c accordingly....use a calculator if you have to

OpenStudy (anonymous):

oh no it would be 12.17

OpenStudy (anonymous):

?